OpenStudy (anonymous):

write the standard form of the equation of the circle with the given center (-2,-4) and radius 6

OpenStudy (anonymous):

using the standard circle equation (x - h)^2 + (y - k)^2 = r^2, where (h,k) is the center. just sub the values in (x + 2)^2 + (y + 4)^2 = 6^2